# HaluMem: Evaluating Hallucinations in Memory Systems of Agents

### v1

@@ 6.2.2 Performance on Different Memory Types
Table 4 reports the extraction accuracy of each memory system for event, persona, and relationship memories, which include all memory points from both the memory extraction and update tasks while excluding distractor memories. Mem0 and Mem0-Graph achieve the best performance on HaluMem-Medium, whereas Supermemory performs best on HaluMem-Long. However, Mem0, Mem0-Graph, and Memobase show a marked decline in long-context scenarios, suggesting difficulty in consistently capturing valuable information in complex dialogues. Supermemory is the only one whose performance on HaluMem-Long surpasses that on HaluMem-Medium, probably because it extracts a larger number of memory points in the long-context condition. Across memory types, Persona memories yield slightly higher accuracy, indicating that static personal traits are easier to capture, whereas understanding event dynamics and relationship changes remains challenging. Overall, all systems still show low performance across the three memory categories, indicating significant limitations in current memory modeling.
@@ 6.2.3 Performance on Different Question Types
Figure 5 illustrates the performance of different memory systems across six categories of questions. Overall, the accuracy of all memory systems remains relatively low across most categories, indicating substantial room for improvement. Except for SuperMemory, the other systems show significantly poorer performance on HaluMem-Long compared to HaluMem-Medium, suggesting a notable degradation under ultra-long context conditions. Among them, Mem0 and Mem0-Graph exhibit the most severe performance drops. In contrast, SuperMemory demonstrates relatively stable behavior and achieves consistently superior overall performance on both datasets. Furthermore, all memory systems perform comparatively better on memory boundary and memory conflict questions, indicating their capability to effectively recognize unknown or misleading information and respond correctly. However, their performance deteriorates substantially on multi-hop inference, dynamic update, and generalization & application questions, suggesting that current memory systems still struggle with complex reasoning and preference tracking.
@@ 6.2.4 Efficiency Analysis of Memory Systems
Dataset System Add Dialogue Retrieve Memory Total
@@ A.4 Construction Details of HaluMem‑Long
The ELI5 dataset consists of factual question–answer pairs (e.g., the second QA example), whereas GPT-OSS-120B-Distilled-Reasoning-math contains question–answer pairs involving mathematics (e.g., the third QA example). To further enrich the diversity of irrelevant dialogues, we also sampled factual QA pairs across eight domains using GPT-4o (e.g., the first example), including Historical Figure, Scientific Concept, Country or Place, Famous Invention, Philosophical Theory, Artwork or Painting, Historical Event, and Mathematical Theorem. These QA pairs are used to simulate dialogues between users and the AI driven by instrumental needs in realistic scenarios. They have minimal impact on the user’s original conversations and do not affect the memory system’s personalized memories of the user. See Appendix D.4 for examples of irrelevant dialogues.
Appendix B Annotation Guidelines and Instructions
B.1 Annotation Objective
@@ B.2 Information Fields
B.3 Annotation Dimensions and Scoring
@@ Appendix C Prompts
C.1 Prompts for Memory Question Answering Task
Figures 7 $\sim$ 9 show the prompt templates used by all memory systems in memory question answering task to assemble questions and retrieve memory points, which are then fed into GPT-4o to generate responses. All of these memory templates are obtained from the official GitHub repositories of the respective memory systems.
@@ C.1 Prompts for Memory Question Answering Task
Figure 9: Prompt for Memobase
C.2 Prompts for Scoring in Memory Evaluation Tasks
Figures 10 $\sim$ 16 respectively illustrate the prompt templates used to guide GPT-4o in scoring for memory extraction, memory updating, and memory question answering tasks.
@@ C.2 Prompts for Scoring in Memory Evaluation Tasks
Appendix D Examples from the Process of Constructing HaluMem
D.1 User Profile Example in Stage 1
@@
D.2 Event Structure Examples in Stage 3
@@
D.3 Examples of Memory Points, Dialogues, and QA Pairs in Stages 4–6
@@
D.4 Examples of irrelevant dialogues

### v2

@@ 6.2.2 Performance on Different Memory Types
Zep 35.76% 39.07% 31.16%
  • * The memory entries of Zep include only those from the memory updating task. For details, see Appendix B.
Table 4 reports the extraction accuracy of each memory system for event, persona, and relationship memories, which include all memory points from both the memory extraction and updating tasks while excluding distractor memories. Zep achieve the best performance on HaluMem-Medium, whereas Supermemory performs best on HaluMem-Long. However, Mem0, Mem0-Graph, and Memobase show a marked decline in long-context scenarios, suggesting difficulty in consistently capturing valuable information in complex dialogues. Supermemory is the only one whose performance on HaluMem-Long surpasses that on HaluMem-Medium, probably because it extracts a larger number of memory points in the long-context condition. Across memory types, Persona memories yield slightly higher accuracy, indicating that static personal traits are easier to capture, whereas understanding event dynamics and relationship changes remains challenging. Overall, all systems still show low performance across the three memory categories, indicating significant limitations in current memory modeling.
@@ 6.2.3 Performance on Different Question Types
Figure 5 illustrates the performance of different memory systems across six categories of questions. Overall, the accuracy of all memory systems remains relatively low across most categories, indicating substantial room for improvement. The Mem0 series and Memobase show significantly poorer performance on HaluMem-Long compared to HaluMem-Medium, suggesting a notable degradation under ultra-long context conditions. In contrast, SuperMemory and Zep demonstrate relatively stable behavior and achieve consistently superior overall performance on both datasets. Furthermore, all memory systems perform comparatively better on memory boundary and memory conflict questions, indicating their capability to effectively recognize unknown or misleading information and respond correctly. However, their performance deteriorates substantially on multi-hop inference, dynamic update, and generalization & application questions, suggesting that current memory systems still struggle with complex reasoning and preference tracking.
@@ 6.2.4 Efficiency Analysis of Memory Systems
Dataset System Dialogue Addition Memory Retrieval Total
@@ A.4 Construction Details of HaluMem‑Long
The ELI5 dataset consists of factual question–answer pairs (e.g., the second QA example), whereas GPT-OSS-120B-Distilled-Reasoning-math contains question–answer pairs involving mathematics (e.g., the third QA example). To further enrich the diversity of irrelevant dialogues, we also sampled factual QA pairs across eight domains using GPT-4o (e.g., the first example), including Historical Figure, Scientific Concept, Country or Place, Famous Invention, Philosophical Theory, Artwork or Painting, Historical Event, and Mathematical Theorem. These QA pairs are used to simulate dialogues between users and the AI driven by instrumental needs in realistic scenarios. They have minimal impact on the user’s original conversations and do not affect the memory system’s personalized memories of the user. See Appendix E.4 for examples of irrelevant dialogues.
Appendix B Special Configurations for Some Memory Systems
This appendix documents the special configurations applied to several memory systems evaluated on HaluMem. While the experimental setup strives to maintain consistent configurations across all evaluated systems, certain memory systems exhibit unique API constraints that necessitate specific adjustments or workarounds. Each subsection below outlines these system-specific configurations to ensure reproducibility.
B.1 Memobase
B.2 Zep
According to our current understanding, the official APIs provided by Zep do not support retrieving all memory points within a specific session, meaning they do not offer functionality equivalent to a Get Dialogue Memory API. Consequently, we were unable to evaluate Zep’s performance on the memory extraction task. We attempted to use the function ‘thread.get_user_context()‘ offered by Zep to obtain all memories under a given thread; however, this method only returns recent memories rather than the complete set, which does not meet the evaluation requirements. Moreover, since Zep’s memory processing workflow operates entirely asynchronously, we could not accurately measure the time consumption in the dialogue addition phase and instead recorded only the time cost associated with memory retrieval.
Appendix C Annotation Guidelines and Instructions
C.1 Annotation Objective
@@ C.2 Information Fields
C.3 Annotation Dimensions and Scoring
@@ Appendix D Prompts
D.1 Prompts for Memory Question Answering Task
Figures 7 $\sim$ 10 show the prompt templates used by all memory systems in memory question answering task to assemble questions and retrieve memory points, which are then fed into GPT-4o to generate responses. All of these memory templates are obtained from the official GitHub repositories of the respective memory systems.
@@ D.1 Prompts for Memory Question Answering Task
You are an intelligent memory assistant tasked with retrieving accurate information from conversation memories. # CONTEXT: You have access to memories from a conversation. These memories contain timestamped information that may be relevant to answering the question. # INSTRUCTIONS: 1. Carefully analyze all provided memories 2. Pay special attention to the timestamps to determine the answer 3. If the question asks about a specific event or fact, look for direct evidence in the memories 4. If the memories contain contradictory information, prioritize the most recent memory 5. If there is a question about time references (like "last year", "two months ago", etc.), calculate the actual date based on the memory timestamp. For example, if a memory from 4 May 2022 mentions "went to India last year," then the trip occurred in 2021. 6. Always convert relative time references to specific dates, months, or years. For example, convert "last year" to "2022" or "two months ago" to "March 2023" based on the memory timestamp. Ignore the reference while answering the question. 7. Focus only on the content of the memories. Do not confuse character names mentioned in memories with the actual users who created those memories. 8. The answer should be less than 5-6 words. # APPROACH (Think step by step): 1. First, examine all memories that contain information related to the question 2. Examine the timestamps and content of these memories carefully 3. Look for explicit mentions of dates, times, locations, or events that answer the question 4. If the answer requires calculation (e.g., converting relative time references), show your work 5. Formulate a precise, concise answer based solely on the evidence in the memories 6. Double-check that your answer directly addresses the question asked 7. Ensure your final answer is specific and avoids vague time references {context} Question: {question} Answer:
Figure 10: Prompt for Zep
D.2 Prompts for Scoring in Memory Evaluation Tasks
Figures 11 $\sim$ 17 respectively illustrate the prompt templates used to guide GPT-4o in scoring for memory extraction, memory updating, and memory question answering tasks.
@@ D.2 Prompts for Scoring in Memory Evaluation Tasks
Appendix E Examples from the Process of Constructing HaluMem
E.1 User Profile Example in Stage 1
@@
E.2 Event Structure Examples in Stage 3
@@
E.3 Examples of Memory Points, Dialogues, and QA Pairs in Stages 4–6
@@
E.4 Examples of irrelevant dialogues